Thanks to Jake, Rock, Zoe, Gary, and the other people who brought HMI back!
Wednesday, April 24, 2013
HMI is Back Online
HMI has returned to normal operations. The HMI team at LMSAL and Stanford will closely monitor the instrument to verify its performance. We don't anticipate any more outages. Data is flowing to the JSOC but it will take a while to prime the pump and produce magnetograms and Dopplergrams.
HMI is Down, WIll Return Later Today
Last night the HMI instrument rebooted and stopped producing science data. The HMI team has made sure the instrument is safe by rebooting and testing HMI. They will restart the science sequencer later today. HMI data will not be available until the sequences are restarted.
Many thanks to the LMSAL and Stanford people who worked through the night to reboot and test HMI!
Wednesday, April 17, 2013
Planned science calibration activities today
Today, starting at 1315 UTC (9:15 am ET), SDO executed the EVE Field
of View maneuver followed by the HMI/AIA Flatfield at 1630 UTC (10:30
am ET). During these maneuvers the science experienced interruption, and some of the browse images reflect the motion of the spacecraft.
Tuesday, April 16, 2013
Brief data stream outage on April 16
We are experiencing a system outage that is interrupting the automatic web updating of SDO images. The spacecraft is still fine and is collecting data as usual, and no SDO data has been lost. The outage is on the ground and is only temporary. We apologize for any inconvenience, and we thank you for your patience.
Update: The image processing queue is catching up, we should be back to normal in a couple of hours. (2:00PM EDT 16 April)
Update #2: Everything is back to normal. Again, sorry for the delay. (4:40pm EDT 16 April)
Update: The image processing queue is catching up, we should be back to normal in a couple of hours. (2:00PM EDT 16 April)
Update #2: Everything is back to normal. Again, sorry for the delay. (4:40pm EDT 16 April)
Thursday, April 11, 2013
Whoa, A Solar Flare!
After weeks of little activity, the Sun erupted with an M6.5 class flare today. At 0711 UTC (3:11 am ET) active region 11719 erupted. AR 11719 is a large area of magnetic field located at 10°N. Another small flare followed as the M6 decayed.
Wednesday, April 10, 2013
EVE Cruciform, April 10, 2013
Today at 1800 UTC (2 pm ET) SDO will execute an EVE cruciform maneuver. SDO will nod back and forth and up and down. These motions allow EVE to calibrate their spectral irradiances.
Wednesday, April 3, 2013
HMI Roll Maneuver, April 3, 2013
Today we ran an HMI roll maneuver from 0550-1230 UTC (00:40-7:30 ET). The entire SDO spacecraft is spun around so that HMI can verify its operation and measure some calibration data.
Subscribe to:
Posts (Atom)
